# Lectures on Gauge theories and Many-Body systems

### v1

@@
Lectures on Gauge theories and Many-Body systems
@@ 1 Introduction
These lectures discuss two correspondences between gauge theories and integrable many-body systems. The first correspondence goes back to the work of many mathematicians and physicists in the 1980-1990’s and is provided by an infinite dimensional Hamiltonian reduction and its quantum counterparts. In this approach the quantization parameters of the gauge theory coincide with the quantization parameters of the many-body system. The second correspondence started emerging in the mid-1990’s, it involves non-trivial dualities, relating classical problems on one side to quantum on another and vice versa. This duality has various reincarnations: Fourier and Legendre transforms, Langlands duality, etc. The quantization parameters are mapped to geometric parameters. Simple questions on one side solve complicated ones on the other and vice versa.
@@
Why many-body systems? Gauge theories describe dynamics of gauge equivalence classes of connections, the simplest observables to define in the classical gauge-theoretic setting are the conjugacy classes of the holonomies of the dynamical gauge field along a fixed loop in space time. If this loop is moved from one location to another the conjugacy class of the holonomy changes. For classical gauge groups these conjugacy classes are in one-to-one correspondence with configurations of points on a circle modulo permutations and, possibly, other symmetries, thus the abovementioned change in the conjugacy class becomes a dynamics of indistinguishable point particles on a circle.
@@
In quantum context the positions of these particles becomes random. Depending on the dimensionality of spacetime, and the degree of supersymmetry of the underlying gauge theory, the mapping of spacetime gauge dynamics to that of a many-body system could be exact or approximate.
@@
In what follows we study the Calogero-Moser-Sutherland class of many-body systems, associated with $A$ type root system and $SU(N)$ gauge theories in various spacetime dimensions, starting in $0+1$ going all the way to $5+1$, with up to eight supercharges. The correspondence will be direct in $0+1,1+1,2+1,3+1$ dimensions for non-supersymmetric matrix quantum mechanics, two dimensional Yang-Mills theory, modified three dimensional Chern-Simons theory, or a hybrid 3d version of 2d Yang-Mills theory, and the hybrid 4d Chern-Simons theory. The correspondence will be of second type for $3+1,4+1,5+1$ dimensional super-Yang-Mills theories with eight supercharges, subject to the so-called $\Omega$-deformation.
@@
We shall introduce a class of non-local observables, generalizing the loop operator of the old-fashioned matrix models. Specifically for supersymmetric gauge theories the correlation functions of these observables obey a non-perturbative version of the Dyson-Schwinger equations that we review. From these equations, in the $3+1$ dimensional case follow the Schrödinger equation of a many-body system. The correspondence in the $4+1$ and $5+1$ dimensional cases remains a conjecture.
@@
IC is grateful to Andrei Okounkov and Aleksandr Artemev for profitable discussions and Anton Dzhamay for his stimulating interest in the project.
@@ 2.1 Definition of the system
Like any classical Hamiltonian system, the Calogero-Moser (CM) system ([Calogero1971]) is a symplectic manifold $(\mathcal{P},{\omega})$ called the phase space and a function $H$ on $\mathcal{P}$, the Hamiltonian.
@@ 2.2 Integrals of motions
$\begin{dcases}\dot{x}_{i}=\frac{\partial H}{\partial p_{i}}=p_{i}\\ \dot{p}_{i}=-\frac{\partial H}{\partial x_{i}}=2\nu^{2}\sum\limits_{k\neq i}\frac{1}{(x_{i}-x_{k})^{3}}.\end{dcases}$
One can easily observe that the quantity $P=\sum\limits_{i=1}^{N}p_{i}$ (total momentum) is constant throughout this dynamics. There are in fact $N$ algebraically independent integrals of motion that are Poisson commuting. When such exist the system is called integrable. Anyway the rest of N2N-2 besides $H,P$ are harder to identify. A convenient way of exhibiting integrability is to find the Lax pair $(L,A)$.
@@
Consider two matrices $L=(L_{ij})_{i,j=1}^{N}$, $A=(A_{i,j})_{i,j=1}^{N}\in\mathrm{Mat}_{\mathbb{C}}(N)$
$L_{ij}$ $=p_{i}\delta_{ij}+(1-\delta_{ij})\frac{\sqrt{-1}\nu}{x_{i}-x_{j}},$ (2.1)
$A_{ij}$ $=(1-\delta_{ij})\frac{\sqrt{-1}\nu}{(x_{i}-x_{j})^{2}}-\delta_{ij}\sum\limits_{k\neq i}\frac{\sqrt{-1}\nu}{(x_{i}-x_{k})^{2}}.$ (2.2)
Prove that the equations of motion (2.1) are equivalent to the Lax equation $\dot{L}=[A,L]$.
From the Lax representation we immediately conclude that eigenvalues of $L$ are the integrals of motions. They have a clear mechanical interpretation of asymptotic momenta $p_{1}^{\pm},...,p_{N}^{\pm}$ for $t\rightarrow\pm\infty$.
@@
The functions $H_{k}$ for k=1,,Nk=1,...,N are obviously algebraically independent. Direct check of Poisson commutativity of $H_{k}$’s is quite difficult. In the following we introduce a larger (in the sense of degrees of freedom) but simpler (in the sense of finding the solutions of EOMs) system. This will lead us to prove the Poisson commutativity of $H_{k}$’s.
@@ 2.3 A simple system
    $T^{*}u(N)=\big\{\big(\sqrt{-1}P,\sqrt{-1}X\big):P,X\in\mathrm{Mat}_{\mathbb{C}}(N),P^{\dagger}=P,X^{\dagger}=X\big\}$
    $T^{*}\mathbb{R}^{N}=\big\{z=v+\sqrt{-1}w:v,w\in\mathbb{R}^{N}\big\}$
@@
    $\omega_{L}=\mathrm{Tr}\big(dP\wedge dX\big)-\sqrt{-1}dz^{\dagger}\wedge dz.$
@@
$\begin{dcases}\partial_{t_{k}}X_{ij}=\frac{\partial\varkappa_{k}}{\partial P_{ij}}=P_{ij}^{k-1}\\ \partial_{t_{k}}P_{ij}=-\frac{\partial\varkappa_{k}}{\partial X_{ij}}=0\\ \partial_{t_{k}}z_{i}=\sqrt{-1}\frac{\partial\varkappa}{\partial\overline{z}_{i}}=0,\end{dcases}$
@@ 2.4 Hamiltonian action
Some symplectic manifolds have a certain kind of symmetry called Hamiltonian action.
@@
    for every $\xi\in\mathrm{Lie}(G)$ the corresponding vector field $V_{\xi}$ is the Hamiltonian vector field $V_{\xi}=v_{H_{\xi}}$:
    $dH_{\xi}=-\omega(V_{\xi},\cdot)$
@@
    the $G$-equivariance condition: for $\xi_{1},\xi_{2}\in\mathrm{Lie}(G)$ $\{H_{\xi_{1}},H_{\xi_{2}}\}=H_{[\xi_{1},\xi_{2}]}$.
Every Hamiltonian action on a symplectic manifold preserves the symplectic form, but there are non-Hamiltonian automorphisms of symplectic manifold that preserve the symplectic form.
@@
$\mu_{L}(P,X,z)=[P,X]-\sqrt{-1}z\otimes z^{\dagger}$ (2.6)
@@
$H_{\xi}(P,X,z)$ $=\mathrm{Tr}\big(\xi([P,X]-\sqrt{-1}z\otimes z^{\dagger})\big).$
@@
$\mu(P,X,z)=[P,X]+\sqrt{-1}(z\otimes z^{\dagger}-\nu 1_{N})$ (2.7)
@@ 2.5 Symplectic reduction
For the discussion below it is usefull to formulate the following
@@
(Meyer [MEYER1973259], Marsden-Weinstein [MARSDEN1974121], see also [woodward2011]) Let $(M,\omega)$ be a symplectic manifold equipped with a Hamiltonian $G$-action with a moment map $\mu$. Suppose also that $G$ acts properly and freely on $\mu^{-1}(0)$. Then $\mu^{-1}(0)/G$ has a structure of a smooth manifold of dimension $\dim M-2\dim G$ with a unique symplectic form $\omega^{\prime}$ satisfying $i^{*}\omega=\pi^{*}\omega^{\prime}$, where $i:\mu^{-1}(0)\rightarrow M$ and $\pi:\mu^{-1}(0)\rightarrow\mu^{-1}(0)/G$ are the inclusion and the projection respectively.
This allows us to produce a new Hamiltonian system starting from some Hamiltonian system with a Hamiltonian action. The construction providing is called the symplectic reduction.
@@
Let $(M,\omega,H)$ be a Hamiltonian system equipped with a Hamiltonian action of a Lie group $G$ with a moment map $\mu$. The Hamiltonian reduction $(M^{\prime},\omega^{\prime},H^{\prime})=(M,\omega,H)/\mkern-5.0mu/G$ is a Hamiltonian system consisting a triple
@@
    Symplectic structure $\omega^{\prime}$ comes from $\omega$ with help of (2.1).
@@
Let $\{\xi_{i}\}_{i}$ be some basis of $\mathrm{Lie}(G)$. By (2.2) $v_{H}.H_{\xi}=\{H,H_{\xi}\}=0$, hence $v_{H}$ is tangent to $\mu^{-1}(0)^{ss}\subset\mu^{-1}(0)=\cap_{i}H_{\xi_{i}}^{-1}(0)$. To prove the second statement we need to check that for any $u\in T_{\pi(m)}M^{\prime}$ $\omega^{\prime}(\pi_{*}v_{H}\big|_{m},u)=-dH(u)$. Let $\tilde{u}\in T_{m}\mu^{-1}(0)^{ss}$ be any lifting of $u$. Using the notation for $i$ and $\pi$ as in (2.1) we have
@@ 2.6 Hamiltonians of the reduction
$\begin{dcases}z_{i}=\sqrt{\nu},\;i=1,...,N\\ \\ P_{ij}=\frac{\sqrt{-1}\nu}{x_{i}-x_{j}},\;i\neq j,\end{dcases}$
@@
$P_{ij}=p_{i}\delta_{ij}+\frac{\sqrt{-1}\nu(1-\delta_{ij})}{x_{i}-x_{j}}.$ (2.8)
@@
Since $\pi^{*}$ is injective this impies that $H_{k}$’s Poisson-commute as well.
@@ 3.1 Lagrangian description
Let us consider the action functional for the system $(\mathcal{P}_{L},\omega_{L},\varkappa)$ in the first-order formalism. We see that $\omega_{L}=d\alpha$ for $\alpha=\mathrm{Tr}PdX+\sqrt{-1}z^{\dagger}dz$. Thus
$S_{L}=\int\alpha-Hdt=\int\mathrm{Tr}(PdX)+\sqrt{-1}z^{\dagger}dz-\frac{1}{2}\mathrm{Tr}P^{2}dt$
Note that this action is not gauge invariant. One can impose the condition $\mu=0$ by introducing a Lagrange multiplier $\sqrt{-1}A_{t}\in u(N)$.
$S=\int\mathrm{Tr}(PdX)+\sqrt{-1}z^{\dagger}dz-\frac{1}{2}\mathrm{Tr}P^{2}dt+\mathrm{Tr}(\mu(P,X,z)A_{t})dt$ (3.1)

### v2

# Gauge theories and Many-Body systems

@@
Gauge theories and Many-Body systems
@@ 1 Introduction
These lectures discuss two correspondences between gauge theories and integrable many-body systems. The first correspondence, based on infinite-dimensional Hamiltonian reduction and its quantum counterparts, grew out of the study of Calogero–Moser-type systems, two-dimensional Yang–Mills theory, and symplectic reduction in the 1980–1990s; see for example [Cal71, Sut71, Mig75, Mey73, MW74, KKS78]. In this approach the quantization parameters of the gauge theory coincide with the quantization parameters of the many-body system. The second correspondence started emerging in the mid-1990s and later acquired a concrete non-perturbative realization in supersymmetric gauge theory and instanton counting; see [Nek02, Nek15]. It involves non-trivial dualities, relating classical problems on one side to quantum ones on the other and vice versa. This duality has various reincarnations: Fourier and Legendre transforms, Langlands duality, and more. The quantization parameters are mapped to geometric parameters. Simple questions on one side solve complicated ones on the other and vice versa.
@@
Why many-body systems? Gauge theories describe dynamics of gauge equivalence classes of connections, and the simplest observables to define in the classical gauge-theoretic setting are the conjugacy classes of the holonomies of the dynamical gauge field along a fixed loop in space time. If this loop is moved from one location to another, the conjugacy class of the holonomy changes. For classical gauge groups these conjugacy classes are in one-to-one correspondence with configurations of points on a circle modulo permutations and, possibly, other symmetries; thus the change in conjugacy class described above becomes the dynamics of indistinguishable point particles on a circle.
@@
In the quantum context, the positions of these particles become random variables. Depending on the dimensionality of spacetime and the degree of supersymmetry of the underlying gauge theory, the mapping of spacetime gauge dynamics to that of a many-body system could be exact or approximate.
@@
In what follows we study the Calogero–Moser–Sutherland class of many-body systems, associated with $A$ type root system and $SU(N)$ gauge theories in various spacetime dimensions, starting in $0+1$ and going all the way to $5+1$, with up to eight supercharges. The correspondence will be direct in $0+1,1+1,2+1,3+1$ dimensions for non-supersymmetric matrix quantum mechanics, two-dimensional Yang–Mills theory, modified three-dimensional Chern–Simons theory, or a hybrid 3d version of 2d Yang–Mills theory, and the hybrid 4d Chern–Simons theory. The correspondence will be of the second type for $3+1,4+1,5+1$ dimensional super-Yang–Mills theories with eight supercharges, subject to the so-called $\Omega$-deformation.
@@
We shall introduce a class of non-local observables, generalizing the loop operator of the old-fashioned matrix models. Specifically for supersymmetric gauge theories the correlation functions of these observables obey a non-perturbative version of the Dyson–Schwinger equations that we review. From these equations, in the $3+1$ dimensional case, follows the Schrödinger equation of a many-body system. The correspondence in the $4+1$ and $5+1$ dimensional cases remains a conjecture.
The first part of these notes develops the symplectic-reduction construction of rational, trigonometric, and elliptic Calogero–Moser systems from gauge-theoretic data. The second part turns to supersymmetric gauge theories, equivariant localization on instanton moduli spaces, measures on partitions, and order/disorder observables. This split mirrors the two correspondences described above.
@@
IC is grateful to Andrei Okounkov and Aleksandr Artemev for useful discussions.
Both of us thank Anton Dzhamay for his stimulating interest in the project.
@@ 2 Calogero–Moser many-body system
The first part of the notes explains how Calogero–Moser-type systems arise from Hamiltonian group actions, moment maps, and symplectic reduction. We begin with the rational model, which already contains the main ideas in their simplest form: a Lax representation, a larger linear Hamiltonian system, and a reduction by a unitary group action. The trigonometric and elliptic systems will then appear as natural variations of this construction and as gauge-theoretic analogues of the same mechanism.
@@ 2.1 Definition of the system
Like any classical Hamiltonian system, the Calogero–Moser (CM) system ([Cal71]) is a symplectic manifold $(\mathcal{P},{\omega})$ called the phase space and a function $H$ on $\mathcal{P}$, the Hamiltonian.
@@ 2.2 Integrals of motions
$\begin{dcases}\dot{x}_{i}=\frac{\partial H}{\partial p_{i}}=p_{i}\\ \dot{p}_{i}=-\frac{\partial H}{\partial x_{i}}=2\nu^{2}\sum\limits_{k\neq i}\frac{1}{(x_{i}-x_{k})^{3}}.\end{dcases}$ (2.1)
One can easily observe that total momentum $P=\sum\limits_{i=1}^{N}p_{i}$ is a constant of motion. There are in fact $N$ algebraically independent integrals of motion that are Poisson commuting. When such exist the system is called integrable. Anyway the rest of N2N-2 besides $H,P$ are harder to identify. A convenient way of exhibiting integrability is to find the Lax pair $(L,A)$.
@@
Consider two matrix-valued functions $L$ and $A$ on $\mathcal{P}$:
$L\,=\,\|L_{ij}({\bf p},{\bf x})\|_{i,j=1}^{N}\,,\ A\,=\,\|A_{i,j}({\bf p},{\bf x})\|_{i,j=1}^{N}\in\mathrm{Mat}_{\mathbb{C}}(N)\,$, (2.2)
where
$L_{ij}({\bf p},{\bf x})$ $=p_{i}\delta_{ij}+(1-\delta_{ij})\frac{{\mathrm{i}}\nu}{x_{i}-x_{j}},$ (2.3)
$A_{ij}({\bf p},{\bf x})$ $=(1-\delta_{ij})\frac{{\mathrm{i}}\nu}{(x_{i}-x_{j})^{2}}-\delta_{ij}\sum\limits_{k\neq i}\frac{{\mathrm{i}}\nu}{(x_{i}-x_{k})^{2}}.$ (2.4)
Prove that the equations of motion (2.1) are equivalent to the Lax equation
$\dot{L}=[A,L]\ .$ (2.5)
From the Lax representation we immediately conclude that eigenvalues of $L$ are the integrals of motions. They have a clear mechanical interpretation of asymptotic momenta
$\left(p_{1}^{\pm},...,p_{N}^{\pm}\right)={\rm Lim}_{t\rightarrow\pm\infty}\left(p_{1}(t)\,,\ldots\,,p_{N}(t)\right)\ .$ (2.6)
@@
The functions $H_{k}$ for k=1,,Nk=1,...,N are algebraically independent. Indeed, in the asymptotic region they become the power sums of the asymptotic momenta $p_{1}^{\pm},\ldots,p_{N}^{\pm}$, and these power sums are algebraically independent on a Zariski-open set (equivalently, their Jacobian is a Vandermonde determinant). A direct check of Poisson commutativity of $H_{k}$’s is quite difficult. In the following we introduce a larger (in the sense of degrees of freedom) but simpler (in the sense of finding the solutions of EOMs) system. This will lead us to a proof of Poisson commutativity of $H_{k}$’s.
@@ 2.3 A simple system
    $T^{*}u(N)=\big\{\big({\mathrm{i}}P,{\mathrm{i}}X\big):P,X\in\mathrm{Mat}_{\mathbb{C}}(N),P^{\dagger}=P,X^{\dagger}=X\big\}$
    $T^{*}\mathbb{R}^{N}=\big\{z=v+{\mathrm{i}}w:v,w\in\mathbb{R}^{N}\big\}$
@@
    $\omega_{L}=\mathrm{Tr}\big(dP\wedge dX\big)-{\mathrm{i}}dz^{\dagger}\wedge dz.$
@@
$\begin{dcases}\partial_{t_{k}}X_{ij}=\frac{\partial\varkappa_{k}}{\partial P_{ij}}=P_{ij}^{k-1}\\ \partial_{t_{k}}P_{ij}=-\frac{\partial\varkappa_{k}}{\partial X_{ij}}=0\\ \partial_{t_{k}}z_{i}={\mathrm{i}}\frac{\partial\varkappa}{\partial\overline{z}_{i}}=0,\end{dcases}$
@@ 2.4 Hamiltonian action
A useful symmetry in the symplectic world is the Hamiltonian group action.
@@
    for every $\xi\in\mathrm{Lie}(G)$, let $V_{\xi}$ denote the vector field generated by the one-parameter subgroup $\exp(t\xi)$. Then $V_{\xi}$ is the Hamiltonian vector field $V_{\xi}=V_{H_{\xi}}$:
    $dH_{\xi}=-\omega(V_{\xi},\cdot)$ (2.8)
@@
    the $G$-equivariance condition: for $\xi_{1},\xi_{2}\in\mathrm{Lie}(G)$
    $\{H_{\xi_{1}},H_{\xi_{2}}\}=H_{[\xi_{1},\xi_{2}]}.$ (2.10)
Every Hamiltonian action on a symplectic manifold preserves the symplectic form. The converse is not true. If $M$ has nontrivial $H^{1}(M,{\mathbb{R}})$, then there are vector fields $V$, which preserve the symplectic form, but do not have a Hamiltonian. Also, if the Lie algebra ${\mathrm{Lie}}(G)$ has nontrivial $H^{2}({\mathrm{Lie}}(G),{\mathbb{R}})$, then even if there exist the Hamiltonians $H_{\xi}$ for each $\xi\in\mathrm{Lie}(G)$ as in (2.8) their Poisson brackets don’t obey (2.10), instead
$\{H_{\xi_{1}},H_{\xi_{2}}\}=H_{[{\xi}_{1},{\xi}_{2}]}+C({\xi}_{1},{\xi}_{2})$ (2.11)
with some constant $C({\xi}_{1},{\xi}_{2})=-({\xi}_{2},{\xi}_{1})$, representing a non-trivial $2$-cocycle on $\mathrm{Lie}(G)$.
@@
$\mu_{L}(P,X,z)=[P,X]-{\mathrm{i}}z\otimes z^{\dagger}$ (2.13)
@@
$H_{\xi}(P,X,z)$ $=\mathrm{Tr}\big(\xi([P,X]-{\mathrm{i}}z\otimes z^{\dagger})\big).$
@@
$\mu(P,X,z)=[P,X]+{\mathrm{i}}(z\otimes z^{\dagger}-\nu 1_{N})$ (2.14)
@@ 2.5 Symplectic reduction
For the discussion below it is useful to formulate the following
@@
(Meyer [Mey73], Marsden-Weinstein [MW74], see also [Woo11]) Let $(M,\omega)$ be a symplectic manifold equipped with a Hamiltonian $G$-action with a moment map $\mu$. Suppose also that $G$ acts properly and freely on $\mu^{-1}(0)$. Then $\mu^{-1}(0)/G$ has a structure of a smooth manifold of dimension $\dim M-2\dim G$ with a unique symplectic form $\omega^{\prime}$ satisfying $i^{*}\omega=\pi^{*}\omega^{\prime}$, where $i:\mu^{-1}(0)\rightarrow M$ and $\pi:\mu^{-1}(0)\rightarrow\mu^{-1}(0)/G$ are the inclusion and the projection respectively.
This allows us to produce a new Hamiltonian system starting from a Hamiltonian system with a Hamiltonian action. The resulting construction is called symplectic reduction.
@@
Let $(M,\omega,H)$ be a Hamiltonian system equipped with a Hamiltonian action of a Lie group $G$ with a moment map $\mu$. The Hamiltonian reduction $(M^{\prime},\omega^{\prime},H^{\prime})=(M,\omega,H)/\mkern-5.0mu/G$ is a Hamiltonian system consisting of a triple
@@
    Symplectic structure $\omega^{\prime}$ is uniquely determined by the identity $i^{*}\omega=\pi^{*}\omega^{\prime}$ from Theorem 2.1.
@@
Let $\{\xi_{i}\}_{i}$ be some basis of $\mathrm{Lie}(G)$. By Remark 2.2 $V_{H}.H_{\xi}=\{H,H_{\xi}\}=0$, hence $V_{H}$ is tangent to $\mu^{-1}(0)^{ss}\subset\mu^{-1}(0)=\cap_{i}H_{\xi_{i}}^{-1}(0)$. To prove the second statement we need to check that for any $u\in T_{\pi(m)}M^{\prime}$ $\omega^{\prime}(\pi_{*}V_{H}\big|_{m},u)=-dH(u)$. Let $\tilde{u}\in T_{m}\mu^{-1}(0)^{ss}$ be any lifting of $u$. Using the notation for $i$ and $\pi$ as in the formulation of the Theorem 2.1 we have
@@ 2.6 Hamiltonians of the reduction
$\begin{dcases}z_{i}=\sqrt{\nu},\;i=1,...,N\\ \\ P_{ij}=\frac{{\mathrm{i}}\nu}{x_{i}-x_{j}},\;i\neq j,\end{dcases}$
@@
$P_{ij}=p_{i}\delta_{ij}+\frac{{\mathrm{i}}\nu(1-\delta_{ij})}{x_{i}-x_{j}}.$ (2.15)
@@
Since $\pi^{*}$ is injective this implies that $H_{k}$’s Poisson-commute as well.
@@ 3.1 Lagrangian description
Let us consider the action functional for the system $(\mathcal{P}_{L},\omega_{L},\varkappa)$ in the first-order formalism. We see that $\omega_{L}=d\alpha$ for $\alpha=\mathrm{Tr}PdX+{\mathrm{i}}z^{\dagger}dz$. Thus
$S_{L}=\int\alpha-Hdt=\int\mathrm{Tr}(PdX)+{\mathrm{i}}z^{\dagger}dz-\frac{1}{2}\mathrm{Tr}P^{2}dt$
Note that this action is not gauge invariant. One can impose the condition $\mu=0$ by introducing a Lagrange multiplier ${\mathrm{i}}A_{t}\in u(N)$.
$S=\int\mathrm{Tr}(PdX)+{\mathrm{i}}z^{\dagger}dz-\frac{1}{2}\mathrm{Tr}P^{2}dt+\mathrm{Tr}(\mu(P,X,z)A_{t})dt$ (3.1)